OpenStudy (anonymous):

Solve the system using elimination. 2x+3y=6 5x-y=-19

OpenStudy (lukecrayonz):

To eliminate one of the variables, make one of equations have equal variables just negative and positive. To do this, multiply the second equation by 3 to get -y to -3y, but you must multiply the whole equation. 15x-3y=-57 2x+3y=6

OpenStudy (lukecrayonz):

Add the equations together now! 15x+2x=17x -57+6=-51 17x=-51 x=-3

OpenStudy (lukecrayonz):

Whoops, made a mistake.

OpenStudy (jiteshmeghwal9):

First equal one of the variable in both of the equations